# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ 6 Conclusions
We have introduced a protocol to perform the teleportation of cat state encoded qubits using beam splitter operations and dispersive measurements. Our protocol is an alternative to the standard CV teleportation protocol, and is preferable when quadrature measurements are not easily performed experimentally. Such systems include those that have an engineerable Jaynes-Cummings Hamiltonian, such as circuit-QED and -QAD systems, and trapped ions, where the types of operations that we have used in the teleportation protocol have been demonstrated in several contexts. At first glance our protocol, summarized in Fig. 2, is reminiscent to standard CV teleportation. However, it differs in the type of measurements that are performed, and the different nature of the states that occur throughout. Parity measurements only partially collapse the wavefunction of a bosonic mode and this results in an additional measurement being required to complete the teleportation. We showed that by using repetitive dispersive measurements it is possible to perform this collapse with high fidelity. Even at the single measurement level, fidelities in the vicinity of $\sim 95\%$ can be attained, which far exceeds the classical bounds for qubit teleportation, at $66\%$ [48].
@@
We introduced two variants of the teleportation protocol, using either two single-mode parity measurements (Variant 1) or a joint parity measurement (Variant 2). Both variants work with similar performance and have a similar dependence with parameters. We also considered parameter choices with real or imaginary choices of $\beta$ and showed that the performance was similar. Joint parity measurements are attractive from the standpoint of the optimality of the number of measurement outcomes. We demonstrated that the displaced dispersive measurement procedure, where the total number of measurements plays a crucial role and can lead to a near-perfect fidelity. Our simulation results show that by choosing the proper optimal interaction time (28) and generating cat states with large absolute eigenvalues $|\alpha|$ and $|\beta|$, to make the coherent states more distinguishable, one can obtain a high level of fidelity. In practice we expect a trade-off when including decoherence as larger amplitude cat states are known to decohere at a faster rate [49, 50]. There should be an optimal choice of amplitude where $|\alpha|,|\beta|$ are large enough that the teleportation protocol works effectively, while the decoherence of the cat states are not excessively large. The nature and strength of the decoherence is rather system dependent and for this reason this is left as future work. However, given that manipulations of cat states are already an experimental reality [43, 36], we believe that our protocol is within the realm of practical implementation.
@@
This work is supported by the National Natural Science Foundation of China (62071301); NYU-ECNU Institute of Physics at NYU Shanghai; the Joint Physics Research Institute Challenge Grant; the Science and Technology Commission of Shanghai Municipality (19XD1423000,22ZR1444600); the NYU Shanghai Boost Fund; the China Foreign Experts Program (G2021013002L); the NYU Shanghai Major-Grants Seed Fund; Tamkeen under the NYU Abu Dhabi Research Institute grant CG008; and the SMEC Scientific Research Innovation Project (2023ZKZD55). MF was supported by the Swiss National Science Foundation Ambizione Grant No. 208886, and by The Branco Weiss Fellowship – Society in Science, administered by the ETH Zürich.
@@ A.1 $\alpha,\beta\in\mathbb{R}$, $\tau=\pi/(2\bar{n})$, $N=1$ measurements
In Sec. 4.3, we verified that $\tau=\pi/(2\bar{n})$ gives a good estimate of the optimal time. By analyzing how the state distribution varies with different parameters, we can gain a more intuitive understanding of the factors that affect the fidelity. In order to check that reasonable projections are being performed to distinguish between the |±$|\pm\chi\rangle$ and |±$|\pm\bar{\chi}\rangle$ states (see Fig. 3), it is illuminating to plot the Fock basis distributions.
@@
Figure 10: The probability distributions in the Fock basis $p_{n}=\langle n|\rho_{\sigma_{a}\sigma_{b}}|n\rangle$ for the single-mode parity measured states $\rho_{\sigma_{a}\sigma_{b}}=\text{Tr}_{b,c}D(\chi)|\psi_{\sigma_{a}\sigma_{b}}\rangle\langle\psi_{\sigma_{a}\sigma_{b}}|D^{\dagger}(\chi)/p_{\sigma_{a}\sigma_{b}}$. Parameters are chosen such that (a) $\beta\in\mathbb{R}$; (b) $\beta\in\mathbb{I}$. The inset shows detail of the $\beta=3i$ case. The dotted lines in the inset show the averaged curve. The dashed lines are $\cos(n\tau)$ for different values of $\beta$ respectively, using $\tau=\pi/(2\bar{n})$. The lines appear as nearly vertical because the scale of the plot is much smaller than the full range of the cosine function $[-1,1]$. Common parameters are $\mu=1/2,\nu=\sqrt{3}/2,\alpha=4,N=1,k_{+}=1,k_{-}=0,\tau=\pi/(2\bar{n})$.
@@
The Fock basis distributions of the states $D(\chi)|\psi_{\sigma_{a}\sigma_{b}}\rangle$ on mode $a$ for various choices of $\alpha,\beta\in\mathbb{R}$ are plotted in Fig. 10(a). We show the probabilities in the region of the |±$|\pm\bar{\chi}\rangle$ peak (see Fig. 3(f)). The peaks |±$|\pm\chi\rangle$ are out of range of the plots and are not visible. We note that the states on mode $a$ are now in a superposition of the |±$|\pm\bar{\chi}\rangle$ states, hence the actual probabilities are not the ideal Poissonian distributions as plotted in Fig. 3(f). For the various values of $\beta$, we see that the zero of the cosine function coincides with the peak of the probabilities. We thus see that $\tau=\pi/(2\bar{n})$ is indeed a nearly optimal choice for removing the distribution. The peak of the distributions shifts to the left with decreasing $|\beta|$ and also broadens. The broadening reduces the effectiveness of the cosine function to remove this peak, leading a relatively lower fidelity. This can be seen in Fig. 11(a), where the superior fidelities are obtained near the top right corner of the plot. We also observe that this plot exhibits a symemtry along the diagonal line $\beta=\alpha$ as both the interaction time (28) and the distribution $p_{n}$ themselves possess a symmetry between $\alpha$ and $\beta$. For the single measurement case $N=1$ with outcome $\sigma_{a}^{\prime}=+$, we find a good fidelity of $\sim 94\%$ can be obtained (see Fig. 11(a)). We expect higher fidelities with larger $|\alpha|$ and $|\beta|$.
@@
Figure 11: Fidelity (43) in the space of $\alpha$ and $\beta$ for Variant 1 teleportation with single-mode parity measurements. We choose parameters such that (a) $\alpha$,$\beta\in\mathbb{R}$; (b) $\alpha\in\mathbb{R}$ and $\beta\in\mathbb{I}$. Common parameters are $\mu=1/2,\nu=\sqrt{3}/2,N=1,k_{+}=1,k_{-}=0,\tau=\pi/(2\bar{n})$.
@@ A.2 $\alpha\in\mathbb{R}$ and $\beta\in\mathbb{I}$ , $\tau=\pi/(2\bar{n})$, $N=1$ measurements
We now consider the equivalent case as Sec. A.1, but with $\beta\in\mathbb{I}$. The Fock basis distributions on mode $a$ of the state $D(\chi)|\psi_{\sigma_{a}\sigma_{b}}\rangle$ are shown in Fig. 10(b). Due to the superposition of the |±$|\pm\bar{\chi}\rangle$ states, the distribution has some interference effects, giving rise to oscillatory behavior (see inset of Fig. 10(b) for further detail). Further, unlike Fig. 10(a), here the $|+\chi\rangle$ state lies closer to the |±$|\pm\bar{\chi}\rangle$ states, resulting in a significant overlap. This can be also seen in Fig. 3(e), where the $|+\chi\rangle$ peak is closer to the |±$|\pm\bar{\chi}\rangle$ peak. This potentially impacts the fidelity especially for small values of $|\alpha|,|\beta|$. This behavior accounts for the sensitive optimal points seen in Fig. 5(b). The interference effects and the presence of the $|+\chi\rangle$ peak can shift the distribution peak, which changes the optimal interaction time. When $\beta=4i$, the ability to distinguish between two states is maximized, and with $\xi\approx 2.06$ with a fidelity level of 95% is reached, as shown in Fig. 5(b). The fidelities for various $\alpha,\beta$ are shown in Fig. 11(b). We first notice that this is also a symmetric pattern, with the better fidelity areas in the top right corner. Moreover, the fidelity are generally slightly lower than the $\beta\in\mathbb{R}$. Despite this, the highest fidelity in fact exceeds the $\beta\in\mathbb{R}$ case, and excellent fidelities are obtained particularly for $|\beta|=|\alpha|$. Again, we expect further improvements of the fidelity with larger $|\alpha|,|\beta|$.
Figure 12: The probability distributions in the Fock basis $p_{n}=\langle n|\rho_{\sigma_{ab}}|n\rangle$ for the joint parity measured states $\rho_{\sigma_{ab}}=\text{Tr}_{b,c}D(\chi)|\psi_{\sigma_{ab}}\rangle\langle\psi_{\sigma_{ab}}|D^{\dagger}(\chi)/p_{\sigma_{ab}}$. Parameters are chosen such that (a) $\beta\in\mathbb{R}$; (b) $\beta\in\mathbb{I}$. The dashed lines are $\cos(n\tau)$ for different values of $\beta$ respectively, using $\tau=\pi/(2\bar{n})$. The lines appear as nearly vertical because the scale of the plot is much smaller than the full range of the cosine function $[-1,1]$. Common parameters are $\mu=1/2,\nu=\sqrt{3}/2,\alpha=4,N=1,k_{+}=1,k_{-}=0,\tau=\pi/(2\bar{n})$.
@@ B.1 $\alpha,\beta\in\mathbb{R}$, $\tau=\pi/(2\bar{n})$, $N=1$ measurements
Figure 13: Fidelity (48) in the space of $\alpha\in\mathbb{R}$ and $\beta\in\mathbb{I}$ for Variant 2 teleportation with joint parity measurements. Parameters are (a) $\alpha$,$\beta\in\mathbb{R}$; (b) $\alpha\in\mathbb{R}$ and $\beta\in\mathbb{I}$. Common parameters are $\mu=1/2,\nu=\sqrt{3}/2,N=1,k_{+}=1,k_{-}=0,\tau=\pi/(2\bar{n})$.

### v2

@@ VII Conclusions
We have introduced a protocol to perform the teleportation of cat state encoded qubits using beam splitter operations and dispersive measurements. Our protocol is an alternative to the standard CV teleportation protocol, and is preferable when quadrature measurements are not easily performed experimentally. Such systems include those that have an engineerable Jaynes-Cummings Hamiltonian, such as circuit-QED and -QAD systems, and trapped ions, where the types of operations that we have used in the teleportation protocol have been demonstrated in several contexts. At first glance our protocol, summarized in Fig. 2, is reminiscent to standard CV teleportation. However, it differs in the type of measurements that are performed, and the different nature of the states that occur throughout. Parity measurements only partially collapse the wavefunction of a bosonic mode and this results in an additional measurement being required to complete the teleportation. We showed that by using repetitive dispersive measurements it is possible to perform this collapse with high fidelity. Even at the single measurement level, fidelities in the vicinity of $\sim 95\%$ can be attained, which far exceeds the classical bounds for qubit teleportation, at $66\%$ Massar and Popescu (1995).
@@
We introduced two variants of the teleportation protocol, using either two single-mode parity measurements (Variant 1) or a joint parity measurement (Variant 2). Both variants work with similar performance and have a similar dependence with parameters. We also considered parameter choices with real or imaginary choices of $\beta$ and showed that the performance was similar. Joint parity measurements are attractive from the standpoint of the optimality of the number of measurement outcomes. We demonstrated that the displaced dispersive measurement procedure, where the total number of measurements plays a crucial role and can lead to a near-perfect fidelity. Our simulation results show that by choosing the proper optimal interaction time (28) and the absolute eigenvalues $|\alpha|,|\beta|$, one can obtain a high level of fidelity. In terms of the protocol itself, a larger $|\alpha|,|\beta|$ gives better results since it satisfies the approximation (14). When including decoherence, selecting larger amplitude of cat states and performing multiple measurements can obtain near-perfect fidelity, despite larger amplitude cat states being more sensitive to loss Zurek (2003); Byrnes and Ilo-Okeke (2021). Given that manipulations of cat states are already an experimental reality Wang et al. (2016); Bild et al. (2023), and the schemes of stabilization of cat states have been proposed Zapletal et al. (2022); Marquet et al. (2024), we believe that our protocol is within the realm of practical implementation.
Acknowledgements.
@@
TB is supported by the SMEC Scientific Research Innovation Project (2023ZKZD55); the Science and Technology Commission of Shanghai Municipality (22ZR1444600); the NYU Shanghai Boost Fund; the China Foreign Experts Program (G2021013002L); the NYU-ECNU Institute of Physics at NYU Shanghai; the NYU Shanghai Major-Grants Seed Fund; and Tamkeen under the NYU Abu Dhabi Research Institute grant CG008. MF was supported by the Swiss National Science Foundation Ambizione Grant No. 208886, and by The Branco Weiss Fellowship – Society in Science, administered by the ETH Zürich.
@@ A.1 $\alpha,\beta\in\mathbb{R}$, $\tau=\pi/(2\bar{n})$, $N=1$ measurements
In Sec. IV.3, we verified that $\tau=\pi/(2\bar{n})$ gives a good estimate of the optimal time. By analyzing how the state distribution varies with different parameters, we can gain a more intuitive understanding of the factors that affect the fidelity. In order to check that reasonable projections are being performed to distinguish between the $|\pm\chi\rangle$ and $|\pm\bar{\chi}\rangle$ states (see Fig. 3), it is illuminating to plot the Fock basis distributions.
@@
Figure 11: The probability distributions in the Fock basis $p_{n}=\langle n|\rho_{\sigma_{a}\sigma_{b}}|n\rangle$ for the single-mode parity measured states $\rho_{\sigma_{a}\sigma_{b}}=\text{Tr}_{b,c}D(\chi)|\psi_{\sigma_{a}\sigma_{b}}\rangle\langle\psi_{\sigma_{a}\sigma_{b}}|D^{\dagger}(\chi)/p_{\sigma_{a}\sigma_{b}}$. Parameters are chosen such that (a) $\beta\in\mathbb{R}$; (b) $\beta\in\mathbb{I}$. The inset shows detail of the $\beta=3i$ case. The dotted lines in the inset show the averaged curve. The dashed lines are $\cos(n\tau)$ for different values of $\beta$ respectively, using $\tau=\pi/(2\bar{n})$. The lines appear as nearly vertical because the scale of the plot is much smaller than the full range of the cosine function [1,1][-1,1]. Common parameters are $\mu=1/2,\nu=\sqrt{3}/2,\alpha=4,N=1,k_{+}=1,k_{-}=0,\tau=\pi/(2\bar{n})$.
@@
The Fock basis distributions of the states $D(\chi)|\psi_{\sigma_{a}\sigma_{b}}\rangle$ on mode $a$ for various choices of $\alpha,\beta\in\mathbb{R}$ are plotted in Fig. 11(a). We show the probabilities in the region of the $|\pm\bar{\chi}\rangle$ peak (see Fig. 3(f)). The peaks $|\pm\chi\rangle$ are out of range of the plots and are not visible. We note that the states on mode $a$ are now in a superposition of the $|\pm\bar{\chi}\rangle$ states, hence the actual probabilities are not the ideal Poissonian distributions as plotted in Fig. 3(f). For the various values of $\beta$, we see that the zero of the cosine function coincides with the peak of the probabilities. We thus see that $\tau=\pi/(2\bar{n})$ is indeed a nearly optimal choice for removing the distribution. The peak of the distributions shifts to the left with decreasing $|\beta|$ and also broadens. The broadening reduces the effectiveness of the cosine function to remove this peak, leading a relatively lower fidelity. This can be seen in Fig. 12(a), where the superior fidelities are obtained near the top right corner of the plot. We also observe that this plot exhibits a symemtry along the diagonal line $\beta=\alpha$ as both the interaction time (28) and the distribution $p_{n}$ themselves possess a symmetry between $\alpha$ and $\beta$. For the single measurement case $N=1$ with outcome $\sigma_{a}^{\prime}=+$, we find a good fidelity of $\sim 94\%$ can be obtained (see Fig. 12(a)). We expect higher fidelities with larger $|\alpha|$ and $|\beta|$.
@@
Figure 12: Fidelity (43) in the space of $\alpha$ and $\beta$ for Variant 1 teleportation with single-mode parity measurements. We choose parameters such that (a) $\alpha$$\beta\in\mathbb{R}$; (b) $\alpha\in\mathbb{R}$ and $\beta\in\mathbb{I}$. Common parameters are $\mu=1/2,\nu=\sqrt{3}/2,N=1,k_{+}=1,k_{-}=0,\tau=\pi/(2\bar{n})$.
@@ A.2 $\alpha\in\mathbb{R}$ and $\beta\in\mathbb{I}$ , $\tau=\pi/(2\bar{n})$, $N=1$ measurements
We now consider the equivalent case as Sec. A.1, but with $\beta\in\mathbb{I}$. The Fock basis distributions on mode $a$ of the state $D(\chi)|\psi_{\sigma_{a}\sigma_{b}}\rangle$ are shown in Fig. 11(b). Due to the superposition of the $|\pm\bar{\chi}\rangle$ states, the distribution has some interference effects, giving rise to oscillatory behavior (see inset of Fig. 11(b) for further detail). Further, unlike Fig. 11(a), here the $|+\chi\rangle$ state lies closer to the $|\pm\bar{\chi}\rangle$ states, resulting in a significant overlap. This can be also seen in Fig. 3(e), where the $|+\chi\rangle$ peak is closer to the $|\pm\bar{\chi}\rangle$ peak. This potentially impacts the fidelity especially for small values of $|\alpha|,|\beta|$. This behavior accounts for the sensitive optimal points seen in Fig. 5(b). The interference effects and the presence of the $|+\chi\rangle$ peak can shift the distribution peak, which changes the optimal interaction time. When $\beta=4i$, the ability to distinguish between two states is maximized, and with $\xi\approx 2.06$ with a fidelity level of 95% is reached, as shown in Fig. 5(b). The fidelities for various $\alpha,\beta$ are shown in Fig. 12(b). We first notice that this is also a symmetric pattern, with the better fidelity areas in the top right corner. Moreover, the fidelity are generally slightly lower than the $\beta\in\mathbb{R}$. Despite this, the highest fidelity in fact exceeds the $\beta\in\mathbb{R}$ case, and excellent fidelities are obtained particularly for $|\beta|=|\alpha|$. Again, we expect further improvements of the fidelity with larger $|\alpha|,|\beta|$.
Figure 13: The probability distributions in the Fock basis $p_{n}=\langle n|\rho_{\sigma_{ab}}|n\rangle$ for the joint parity measured states $\rho_{\sigma_{ab}}=\text{Tr}_{b,c}D(\chi)|\psi_{\sigma_{ab}}\rangle\langle\psi_{\sigma_{ab}}|D^{\dagger}(\chi)/p_{\sigma_{ab}}$. Parameters are chosen such that (a) $\beta\in\mathbb{R}$; (b) $\beta\in\mathbb{I}$. The dashed lines are $\cos(n\tau)$ for different values of $\beta$ respectively, using $\tau=\pi/(2\bar{n})$. The lines appear as nearly vertical because the scale of the plot is much smaller than the full range of the cosine function [1,1][-1,1]. Common parameters are $\mu=1/2,\nu=\sqrt{3}/2,\alpha=4,N=1,k_{+}=1,k_{-}=0,\tau=\pi/(2\bar{n})$.
@@ B.1 $\alpha,\beta\in\mathbb{R}$, $\tau=\pi/(2\bar{n})$, $N=1$ measurements
Figure 14: Fidelity (48) in the space of $\alpha\in\mathbb{R}$ and $\beta\in\mathbb{I}$ for Variant 2 teleportation with joint parity measurements. Parameters are (a) $\alpha$$\beta\in\mathbb{R}$; (b) $\alpha\in\mathbb{R}$ and $\beta\in\mathbb{I}$. Common parameters are $\mu=1/2,\nu=\sqrt{3}/2,N=1,k_{+}=1,k_{-}=0,\tau=\pi/(2\bar{n})$.